Roll Corporation (RC) currently has 505,000 shares of stock outstanding that sell for $80 per share. Assuming no market imperfections or tax effects exist, what will the share price be after:

a. RC has a four-for-three stock split? (Do not round intermediate calculations and round your answer to 2 decimal places, e.g., 32.16.)

New share price $

b. RC has a 20 percent stock dividend? (Do not round intermediate calculations and round your answer to 2 decimal places, e.g., 32.16.)

New share price $

c. RC has a 42.5 percent stock dividend? (Do not round intermediate calculations and round your answer to 2 decimal places, e.g., 32.16.)

New share price $

d. RC has a three-for-seven reverse stock split? (Do not round intermediate calculations and round your answer to 2 decimal places, e.g., 32.16.)

New share price $ e. Determine the new number of shares outstanding in parts (a) through (d). (Do not round intermediate calculations and round your answer to the nearest whole number, e.g., 32.)

a. New shares outstanding

b. New shares outstanding

c. New shares outstanding

d. New shares outstanding
Answer #1

Answers

a) new share price =$60.00

b) new share price =$80.00

c) new share price =$80.00

d) new share price =$186.67

e) new shares outstanding shares in parts from a through d

a) new share outstanding =673,333

b) New shares outstanding =606,000

c) new shares outstanding =719,625

d) new shares outstanding = 216,429

explanation

a) share split four for three

total outstanding capital =505,000*$8= $40,400,000

number of outstanding shares = 505,000*4/3= 673,333.3333333333

new share price = $40,400,000/673,333.3333333333= $60.00

b) stock dividend 20%

stock dividend = 505,000*20/100= 101,000 shares

total outstanding shares = 505,000+101,000= 606,000

**new share price = $80 ( when stock dividend declared share price remains unchanged )

c) stock dividend 42.5%

stock dividend = 505,000*42.5/100=214,625

total outstanding shares = 505,000+214,625 =719,625

new share price = $80 ( share price remains unchanged when stock dividend declared)

d) three for seven stock split

reverse stock split = 505,000*3/7=216,428.5714285714 outstanding shares  

new share price= $40,400,000/216,428.5714285714= $186.67

1.Stock split is the process of increasing the number of outstanding shares and decreasing the par value of stock without changing the total amount of outstanding share capital

2.reverse stock split is the process of merging or combining the shares, it decreases the total number of outstanding shares and increases the share price without changing the total amount of outstanding share capital.

3. Stock dividend is also increases the number of outstanding shares and total outstanding share capital without affecting the par value of the stock

Note : answers are very sensitive because round off factors given so please use comment box for further clarification.